OpenStudy (anonymous):

The LCM of two numbers is 192 and their GCF is 16. One of the numbers is 48. What is the other number?

OpenStudy (anonymous):

Product of two numbers = GCF * LCM

OpenStudy (anonymous):

Let other number be x therefore 48x = 192 * 16 x = 64 so other number is 64

OpenStudy (anonymous):

true thanks alot *